# Model of incompressible turbulent flows via a kinetic theory

### v1

@@ Appendix A Derivations of the Reynolds stress and turbulent kinetic energy flux via Chapman–Enskog expansion
The above method is valid in a continuous velocity space $\xi_{y}$. In practical computations, the velocity space is divided into a finite set of discrete velocities $\left\{\xi_{y,k}\mid k=1,2,\ldots,N\right\}$, where $N$ is the number of discrete velocities. It is important to discretize the velocity space based on certain quadrature rules, which are chosen according to the rarefaction degree. The gradients of VDFs in the source term are approximated by central differences,
$\left(\frac{\partial\Phi_{\alpha}}{\partial\xi_{y}}\right)_{k}=\frac{\Phi_{\alpha}(\xi_{y,k+1})-\Phi_{\alpha}(\xi_{y,k-1})}{(\xi_{y,k+1}-\xi_{y,k-1})}.$ (100)
The discrete velocity method described above is similar to the well-known discrete unified gas kinetic scheme (guo2013discrete), with the only difference being that the construction of the interface distribution function does not take into account collision effects. For completeness, the general algorithmic framework consists of the following steps:
@@
    Evolve the VDFs $\Phi_{\alpha}$ at each cell interface at time $t^{n+1/2}$ according to (94) and (95), and then calculate the mean variables and $\epsilon$ at each cell interface according to (96);
@@
Appendix C The Reduced distribution function $\Phi_{4}$ from kinetic model
The fourth reduced distribution function $\Phi_{4}$, which represents the vertical Reynolds normal stress component, exhibits a behaviour similar to $\Phi_{1}$, remaining essentially in equilibrium across all wall-normal position.
Figure 12: The reduced velocity distribution function $\Phi_{4}$ from the LR-BGK at $\text{Re}=10133$ as a function of the microscopic velocity $\xi_{y}$ scaled by local parameters $\sqrt{4K/3}$ at different wall-normal positions from top to bottom: (a) $y^{+}=0.48$, (b) $y^{+}=11.6$, (c) $y^{+}=215.5$, (d) $y^{+}=419.8$, and (e) $y^{+}=460.4$. Dash lines represent the computed distribution functions, while solid lines represent the local equilibrium distribution functions.
Figure 13: The reduced velocity distribution function $\Phi_{4}$ from the HR-BGK model at $\text{Re}=10133$ as a function of the microscopic velocity $\xi_{y}$ scaled by local parameters $\sqrt{4K/3}$ at different wall-normal positions from top to bottom: (a) $y^{+}=11.3$, (b) $y^{+}=215.3$, (c) $y^{+}=419.2$, and (d) $y^{+}=453.2$. Dash lines represent the computed distribution functions, while solid lines represent the local equilibrium distribution functions.

### v2

@@ Appendix A Derivations of the Reynolds stress and turbulent kinetic energy flux via Chapman–Enskog expansion
The above method is valid in a continuous velocity space $\xi_{y}$. In practical computations, the velocity space is divided into a finite set of discrete velocities $\left\{\xi_{y,k}\mid k=1,2,\ldots,N\right\}$, where $N$ is the number of discrete velocities. With the discrete velocity space, the moments of a continuous distribution function have to be expressed as discrete moments,
$\int\phi(\xi_{y})F(\xi_{y})d\xi_{y}=\sum_{k}w_{k}\phi(\xi_{y,k})F(\xi_{y,k}),$ (92)
where $\phi$ is a polynomial of the particle velocity $\xi_{y}$ and $\omega_{k}$ is the weight of the numerical quadrature at the discrete velocity $\xi_{y,k}$. It is important to discretize the velocity space based on certain quadrature rules, which are chosen according to the rarefaction degree. The gradients of VDFs in the source term are approximated by central differences.
For completeness, the general algorithmic framework consists of the following steps:
@@
    Evolve the VDFs $\Phi_{\alpha}$ at each cell interface at time $t^{n+1/2}$ according to (86) and (87), and then calculate $\epsilon$ at each cell interface according to (88);